OpenStudy (anonymous):

If f(x) = (x^2-x)/(x^2+2x). Then find derivative of inverse of f(x).

OpenStudy (goformit100):

first find the inverse, and then find its derivative OK....

OpenStudy (anonymous):

inverse is so complicated to find :o

OpenStudy (goformit100):

@Koikkara is an Engg. He will help you better.

hartnn (hartnn):

Let \(\Large f(x)=y \implies f^{-1}f(x)=x= f^{-1}(y)\) so, we need derivative of \( f^{-1}(y)\) or of x, which is \(\huge \dfrac{dx}{dy}\) use this : \(\huge \dfrac{dx}{dy}=\dfrac{1}{\dfrac{dy}{dx}}\) so, you'll need dy/dx which is easy, and just take the reciprocal.

OpenStudy (anonymous):

And I shall differentiate whole equation w.r.t y ? :o How ?

hartnn (hartnn):

no no....you'll first find dy/dx by diff. w.r.t x only. then since you required dx/dy, you will take the reciprocal

hartnn (hartnn):

find f'(x)

OpenStudy (anonymous):

So required derivative is 1/f'(x) ??

hartnn (hartnn):

yes!

OpenStudy (anonymous):

Ah Thanks !!

hartnn (hartnn):

welcome ^_^